Q(1) [10 Marks] [CLO2,C3] (a) You take an exam that contains 20 multiple-choice questions. Each question has 4 possible options. You know the answer to 10 questions, but you have no idea about the other 10 questions so you choose answers randomly. Your score X on the exam is the total number of correct answers. Find the PMF of X. Find P(X > 15). (b) Packets at a certain node on the internet arrive with a rate of 100 packets per minute. Find the probability that no packets arrive in 6 seconds. Find the probability that 2 or more packets arrive in the first 6 seconds.


1
Expert's answer
2021-04-28T06:49:41-0400

Solution (a):

Let's define the random variable Y as the number of your correct answers to the 10 questions you answer randomly. Then your total score will be X=Y+10. First, let's find the PMF of Y. For each question, your success probability is "\frac{1}{4}" . Hence, you perform 10 independent "\operatorname{Bernoulli}\left(\frac{1}{4}\right)" trials and Y is the number of successes. Thus, we conclude "Y \sim Binomial \left(10, \frac{1}{4}\right)" , so

"P_{Y}(y)=\left\{\begin{array}{ll}

\left(\begin{array}{c}

10 \\

y

\end{array}\right)\left(\frac{1}{4}\right)^{y}\left(\frac{3}{4}\right)^{10-y} & \text { for } y=0,1,2,3, \ldots, 10 \\

0 & \text { otherwise }

\end{array}\right."

Now we need to find the PMF of X=Y+10. First note that "R_{X}=\{10,11,12, \ldots, 20\}" . We can write


"\begin{aligned}

P_{X}(10) &=P(X=10)=P(Y+10=10) \\

&=P(Y=0)=\left(\begin{array}{c}

10 \\

0

\end{array}\right)\left(\frac{1}{4}\right)^{0}\left(\frac{3}{4}\right)^{10-0}=\left(\frac{3}{4}\right)^{10} \\

P_{X}(11) &=P(X=11)=P(Y+10=11) \\

&=P(Y=1)=\left(\begin{array}{c}

10 \\

1

\end{array}\right)\left(\frac{1}{4}\right)^{1}\left(\frac{3}{4}\right)^{10-1}=10\left(\frac{1}{4}\right)\left(\frac{3}{4}\right)^{9} .

\end{aligned}"

In general for "k \in R_{X}=\{10,11,12, \ldots, 20\}" ,

"\begin{aligned}

P_{X}(k) &=P(X=k)=P(Y+10=k) \\

&=P(Y=k-10)=\left(\begin{array}{l}

10 \\

k-10

\end{array}\right)\left(\frac{1}{4}\right)^{k-10}\left(\frac{3}{4}\right)^{20-k} 

\end{aligned}"

Thus,

"P_{X}(k)=\left\{\begin{array}{ll}

\left(\begin{array}{l}

10 \\

k-10

\end{array}\right)\left(\frac{1}{4}\right)^{k-10}\left(\frac{3}{4}\right)^{20-k} & \text { for } k=10,11,12, \ldots, 20 \\

0 & \text { otherwise }

\end{array}\right."

In order to calculate P(X>15), we know we should consider y=6,7,8,9,10

"\begin{array}{c}
P_{Y}(y)=\left\{\begin{array}{ll}
\left(\begin{array}{c}
10 \\
y
\end{array}\right)\left(\frac{1}{4}\right)^{y}\left(\frac{3}{4}\right)^{10-y} & \text { for } y=6,7,8,9,10 \\
0 & \text { otherwise }
\end{array}\right. \\
P_{X}(k)=\left\{\begin{array}{ll}
\left(\begin{array}{c}
10 \\
k-10
\end{array}\right)\left(\frac{1}{4}\right)^{k-10}\left(\frac{3}{4}\right)^{20-k} & \text { for } k=16,17, \ldots, 20 \\
0 & \text { otherwise }
\end{array}\right. \\
P(X>15)=P_{X}(16)+P_{X}(17)+P_{X}(18)+P_{X}(19)+P_{X}(20) \\
=\left(\begin{array}{c}
10 \\
6
\end{array}\right)\left(\frac{1}{4}\right)^{6}\left(\frac{3}{4}\right)^{4}+\left(\begin{array}{c}
10 \\
7
\end{array}\right)\left(\frac{1}{4}\right)^{7}\left(\frac{3}{4}\right)^{3}+\left(\begin{array}{c}
10 \\
8
\end{array}\right)\left(\frac{1}{4}\right)^{8}\left(\frac{3}{4}\right)^{2} \\
+\left(\begin{array}{c}
10 \\
9
\end{array}\right)\left(\frac{1}{4}\right)^{9}\left(\frac{3}{4}\right)^{1}+\left(\begin{array}{c}
10 \\
10
\end{array}\right)\left(\frac{1}{4}\right)^{10}\left(\frac{3}{4}\right)^{0}
\end{array}"

"=0.019727"

(b):

Given, "\lambda=100" per minute"=\dfrac{100}{60}\ or\ \dfrac{5}{3}" per second

For 6 second, "\lambda=\dfrac{5}{3}\times 6=10"

"P(X=0)=10^0\dfrac{e^{-10}}{0!}=0.00004539"

Next, "P(X\ge2)=1-P(X<2)=1-[P(X=0+P(X=1)]"

"=1-(0.00004539+10^1\dfrac{e^{-10}}{1!})=0.9995006"